# A conjecture on descents, inversions and the weak order on Coxeter groups

### v2

@@ 4.1. Type $B$ permutations, inversions and standardization
$\begin{array}[]{ccc}\dots\bar{n}&{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\dots|\dots}&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}n\dots}\\ \mathcal{F}(w)&{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathcal{L}(w)}&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{R}(w)},\end{array}$
i.e., it contains first all the forgotten letters ending with $\bar{n}$ all in negative positions, then the left letters in negative then positive positions, then all the right letters starting with $n$. Moreover $\overline{\mathcal{F}(w)}=\mathcal{R}(w)$ and $\mathcal{L}(w)$ is a symmetric subset of $[\![\bar{n},n]\!]^{*}$.
@@
This is direct from the definition. Since $n$ is in position $k>0$ and $\bar{n}$ is in position $\bar{k}$. The right letters are all letters $a<n$ such that $(a,n)$ is an inversion of $w$: they are all the letters placed to the right of $n$, in positive positions by definition. The forgotten letters are the all the opposite of the right letters which by symmetry of $w$ are placed in all negative positions before $\bar{k}$. In the middle of the word, we obtain the set $\mathcal{L}(w)$ of left letter, which is a symmetric subset of $[\![\bar{n},n]\!]^{*}$. ∎
The following proposition is the equivalent in the case of type $B$ tof Propositions 3.11 in type $A$.
@@
Let $w\in W_{n}$ such that $n$ is in positive position in $w$. Let $\{u,v\}$ be a bipartition of an element $w$. Then the left words $\{u_{\mathcal{L}},v_{\mathcal{L}}\}$ form a bipartition of $w_{\mathcal{L}}$ and the right words $\{u_{\mathcal{R}},v_{\mathcal{R}}\}$ form a bipartition of $w_{\mathcal{R}}$ and
(7) $d_{R}(w)$ $=d_{R}(w_{\mathcal{L}})+d_{R}(w_{\mathcal{R}}),$
(8) $d_{R}(u)$ $=d_{R}(u_{\mathcal{L}})+d_{R}(u_{\mathcal{R}}),$
(9) $d_{R}(v)$ $=d_{R}(v_{\mathcal{L}})+d_{R}(v_{\mathcal{R}}).$
Note that $w_{\mathcal{L}}$, $u_{\mathcal{L}}$, $v_{\mathcal{L}}$ and $w_{\mathcal{R}}$, $u_{\mathcal{R}}$, $v_{\mathcal{R}}$ belong to different types of Coxeter groups: the left words are of type $B$ and the right words are of type $A$.
@@
Here, $w=4\bar{5}\bar{2}\bar{6}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}31|\bar{1}\bar{3}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}625\bar{4}}$ with $u=4\bar{5}\bar{2}\bar{6}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}3\bar{1}|1\bar{3}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}625\bar{4}}$ and $v=\bar{6}\bar{5}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bar{4}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bar{3}}\bar{2}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1}|{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bar{1}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}3}4{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}56}$ forming a bipartition of $w$.We obtain that $u_{\mathcal{L}}=2\bar{1}|1\bar{2}$ and $v_{\mathcal{L}}=\bar{2}1|\bar{1}2$ form a bipartition of $w_{L}=21|\bar{1}\bar{2}$ in $W_{2}$, while $u_{\mathcal{R}}=4231$ and $v_{\mathcal{R}}=1234$ form a (trivial) bipartition of $w_{\mathcal{R}}=u_{\mathcal{R}}$ in $\mathcal{S}_{3}$. Moreover, we have
44 44 $=$ $d_{R}(w)$ $=$ $d_{R}(4\bar{5}\bar{2}\bar{6}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}31|.\bar{1}.\bar{3}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6.25.\bar{4}})$ $=$ $d_{R}(21|.\bar{1}.\bar{2})+d_{R}(4.23.1)$ $=$ $2+2$
33 33 $=$ $d_{R}(u)$ $=$ $d_{R}(4\bar{5}\bar{2}\bar{6}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}3\bar{1}|1.\bar{3}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6.25.\bar{4}})$ $=$ $d_{R}(2\bar{1}|1.\bar{2})+d_{R}(4.23.1)$ $=$ $1+2$
11 11 $=$ $d_{R}(v)$ $=$ $d_{R}(\bar{6}\bar{5}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bar{4}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bar{3}}\bar{2}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1|.\bar{1}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}3}4{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}56})$ $=$ $d_{R}(\bar{2}1|.\bar{1}2)+d_{R}(1234)$ $=$ $1+0$
@@
As in type $A$, the proof is by induction on $\mathbb{N}^{*}$. The result is true for $n=1$ since $W_{1}$ is of type $B_{1}=A_{1}$. Now, let $\{u,v\}$ be a bipartition of an element $w\in W_{n}$ with $n\geq 2$. We suppose that the statement of Conjecture 1 is satisfied for all $w\in W_{k}$ with $1\leq k<n$ such that $n$ is in positive position in $w$.
Suppose first that $|\mathcal{L}(w)|=k$ with $k>0$. By Proposition 4.18, $\{u_{\mathcal{L}},v_{\mathcal{L}}\}$ is a bipartition of $w_{\mathcal{L}}$, and $\{u_{\mathcal{R}},v_{\mathcal{R}}\}$ is a bipartition of $w_{\mathcal{R}}$, $w_{\mathcal{L}}$ is an element of $W_{k}$, with $k<n$, and $w_{\mathcal{R}}$ is an element of $A_{n-1-k}$. Therefore $\{u_{\mathcal{L}},v_{\mathcal{L}}\}$ satisfies Conjecture 1 by induction while $\{u_{\mathcal{R}},v_{\mathcal{R}}\}$ satisfies Conjecture 1 using Theorem 1.5 in the case of type $A$. Hence:
$d_{R}(w_{\mathcal{L}})=d_{R}(u_{\mathcal{L}})+d_{R}(v_{\mathcal{L}})\textrm{ and }d_{R}(w_{\mathcal{R}})=d_{R}(u_{\mathcal{R}})+d_{R}(v_{\mathcal{R}}).$
@@
Now suppose that $\mathcal{L}(w)=\emptyset$. Therefore $n$ is positive position $1$ (as in Example 4.16). Hence $w_{\mathcal{R}}\not=w$ is a standardization of the letters in positive positions in $w$ and an element of $\mathcal{S}_{n-1}$ such that $d_{R}(w_{\mathcal{R}})=d_{R}(w)$. By Proposition 4.18, $\{u_{\mathcal{R}},v_{\mathcal{R}}\}$ form a bipartition of $w_{\mathcal{R}}$ with $d_{R}(u_{R})=d_{R}(u)$, $d_{R}(v_{\mathcal{R}})=d_{R}(w)$. Applying Theorem 1.5 in the case of type type $A$, we obtain $d_{R}(w)=d_{R}(u)+d_{R}(v)$. ∎
@@
Let $w\in W_{n}$ such that $n$ is in positive position in $w$. Let $\{u,v\}$ be a bipartition of an element $w$. Then there is no inversion $(a,c)$ of $u$ or $v$ such that $a\in\mathcal{F}(w)\cup\mathcal{L}(w)$ and $c\in\mathcal{R}(w)$, nor such that $a\in\mathcal{F}(w)$ and $c\in\mathcal{L}(w)\cup\mathcal{R}(w)$. In particular, $n$ is placed after all letters of $\mathcal{F}(w)$ and $\mathcal{L}(w)$ in both $u$ and $v$.
@@
Let $w\in W_{n}$ such that $n$ is in positive position in $w$. Let $\{u,v\}$ be a bipartition of an element $w$. If there is a descent in the word $u$ (or in the word $v$) between two letters $c$ and $a$, then either $a,c\in\mathcal{R}(w)$, or $a,c\in\mathcal{L}(w)$, or $a,c\in\mathcal{F}(w)$. In other words, a descent can only happen between two letters of similar colors.
@@
Let $w\in W_{n}$ such that $n$ is in positive position in $w$. Let $\{u,v\}$ be a bipartition of an element $w$. Then if there is a descent in $u$ (or in $v$) between two letters $c$ and $a$, then either $a,c\in\mathcal{R}(w)$ or $a,c\in\mathcal{L}(w)\cup\mathcal{F}(w)$.
@@
Consider the form of the word $w$ in Lemma 4.17 and observe that the definition of the left letters in type $A$ (everything before the letter $n$ in $w$) corresponds to $\mathcal{L}(w)\cup\mathcal{F}(w)$. Similarly, the right letters of type $A$ correspond to $\mathcal{R}(w)$ in type $B$. The exact same line of reasoning used to prove Proposition 3.17 applies here with Lemma 4.20 and Proposition 4.5. ∎
@@
Suppose that $u$ is of the form $\dots ca\dots$ with $a<c$. Note that we do not specify if the descent is in positive, negative or 00 position as this does not impact the proof. In particular, we could have $a=\bar{c}$ if the descent is in position 00. Lemma 4.23 states that either $a,c\in\mathcal{R}(w)$ or $a,c\in\mathcal{L}(w)\cup\mathcal{F}(w)$. By Lemma 4.20, we cannot have $c\in\mathcal{L}(w)$ and $a\in\mathcal{F}(w)$. The only case to consider is then $c\in\mathcal{F}(w)$ and $a\in\mathcal{L}(w)$ which we suppose by contradiction. As $\mathcal{L}(w)$ is symmetric, we also have $\bar{a}\in\mathcal{L}(w)$ and in particular $\bar{a}\neq c$ and the descent is not in position 00. By symmetry, this implies that there is a descent between $\bar{a}\in\mathcal{L}(w)$ and $\bar{c}\in\mathcal{R}(w)$ (because $\mathcal{F}(w)=\overline{\mathcal{R}(w)}$) which is impossible accordingly to Lemma 4.23. ∎
@@
Let $w\in W_{n}$ such that $n$ is in positive position in $w$. Let $\{u,v\}$ be a bipartition of an element $w$. Assume that $u$ (or $v$) is of the form $\dots b\dots c\dots$ with $b$ and $c$ both in the same set $I$ such that: (1) $I=\mathcal{L}(w)$, $I=\mathcal{R}(w)$ or $I=\mathcal{F}(w)$; (2) the subword between $b$ and $c$ is non-empty all its letters are in $[\![\bar{n},n]\!]^{*}\setminus I$. Then $b<c$.
@@
We suppose by contradiction that $u=\dots cu^{\prime}b\dots$ with $b<c$ and $b,c\in I$ with $I=\mathcal{L}(w)$, $I=\mathcal{R}(w)$ or $I=\mathcal{F}(w)$ such that $u^{\prime}$ is not empty and with letters not in $I$. We proceed to a case by case analysis and show that we always reach a contradiction
Case $b,c\in\mathcal{L}(w)$
The word $u^{\prime}$ is non-empty by hypothesis, we call $a$ the last letter of $u^{\prime}$ so $u$ is of the form $\dots c\dots ab\dots$. As $a\notin\mathcal{L}(w)$, there is no descent between $a$ and $b$ by Proposition 4.21 so $a<b$. In particular, $(a,c)\in\operatorname{invs}(u)$. As $c\in\mathcal{L}(w)$, by Lemma 4.20, we cannot have $a\in\mathcal{F}(w)$, so $a\in\mathcal{R}(w)$. Besides, as $(a,c)\in\operatorname{invs}(u)$, we have $(a,c)\not\in\operatorname{invs}(v)$. By Lemma 4.20, we also know that $n$ is placed after all letters of $\mathcal{L}(w)$ in both $u$ and $v$, this gives
@@
We see that $(a,n)\not\in\operatorname{invs}(u)$ and $(a,n)\not\in\operatorname{invs}(v)$. This contradicts the fact that $(a,n)\in\operatorname{invs}(w)$ since $a$ is a right letter of $w$.
Case $b,c\in\mathcal{R}(w)$
We call $a$ the last letter of $u^{\prime}$ so that $u=\dots c\dots ab\dots$. As before, because of Proposition 4.21, $a<b$. We have $a\in\mathcal{L}(w)\cup\mathcal{F}(w)$ and $(a,c)\in\operatorname{invs}(u)$: this contradicts Lemma 4.20.
Case $b,c\in\mathcal{F}(w)$
Obtained by symmetry of the case $b,c\in\mathcal{R}(w)$. ∎
@@
Let $w\in W_{n}$ such that $n$ is in positive position in $w$. Let $\{u,v\}$ be a bipartition of an element $w$. By a similar argument as in the case of type $A$, we obtain that $\{u_{\mathcal{L}},v_{\mathcal{L}}\}$ form a bipartition of $w_{\mathcal{L}}$ and $\{u_{\mathcal{R}},v_{\mathcal{R}}\}$ form a bipartition of $w_{\mathcal{R}}$, by simply restricting the sets of inversions to the left word (resp. right word) letters and using Lemma 4.4.
We first show Equality (7). We use Lemma 4.17 and suppose that $n$ is in position $k>0$. There is a descent at position $10\leq i<k-$ 1 in $w$ if and only if there is a descent at position $i$ in $w_{\mathcal{L}}$ as $w_{\mathcal{L}}$ is the type $B$ standardization of the middle factor of $w$. Similarly, there is a descent at position $i>k$ in $w$ if and only if there is a descent at position $i-k+1$ in $w_{\mathcal{R}}$. Besides, there is no descent in $w$ at position $k-1$. We find $d_{R}(w)=d_{R}(w_{\mathcal{L}})+d_{R}(w_{\mathcal{R}})$ (as $d_{R}(w)$ and $d_{R}(w_{\mathcal{L}})$ only count descents in nonnegative positions).
@@
$d_{R}(u)\leq d_{R}(u_{\mathcal{L}})+d_{R}(u_{\mathcal{R}}).$
Indeed, suppose that we have a descent $ca$ in $u$ in positive position. Either $a,c\in\mathcal{L}(w)$ and the descent is kept in $u_{\mathcal{L}}$, still in positive position. Or $a,c\in\mathcal{R}(w)$ and the descent is kept in $u_{\mathcal{R}}$. Finally, if $a,c\in\mathcal{F}(w)$, then $u$ also has a descent in negative position between $\bar{a}$ and $\bar{c}$ in $\mathcal{R}(w)$ which was not counted in $d_{R}(u)$ but is kept in $u_{\mathcal{R}}$ and counted in $d_{R}(u_{\mathcal{R}})$.

### v3

@@ 4.1. Type $B$-permutations, inversions and standardization
[…]
i.e., it contains first all the forgotten letters ending with $\bar{n}$ all in negative positions, then the left letters in negative then positive positions, then all the right letters starting with $n$. Moreover $\overline{\mathcal{F}(w)}={\overrightarrow{n}}(w)$ and ${\overleftarrow{n}}(w)$ is a symmetric subset of $\bar{n},n]\!]^{*}$.
@@
This is direct from the definition. Since $n$ is in position $k>0$ and $\bar{n}$ is in position $\bar{k}$. The right letters are all letters $a<n$ such that $(a,n)$ is an inversion of $w$: they are all the letters placed to the right of $n$, in positive positions by definition. The forgotten letters are exactly the opposite of the right letters which by symmetry of $w$ are placed in all negative positions before $\bar{k}$. In the middle of the word, we obtain the set ${\overleftarrow{n}}(w)$ of left letter, which is a symmetric subset of $\bar{n},n]\!]^{*}$. ∎
The following proposition is the equivalent in the case of type $B$ of Propositions 3.13 in type $A$.
@@
Let $w\in W_{n}$ such that $n$ is in positive position in $w$. Let $\{u,v\}$ be a bipartition of an element $w$. Then the left words $\{u_{\overleftarrow{n}},v_{\overleftarrow{n}}\}$ form a bipartition of $w_{\overleftarrow{n}}$ and the right words $\{u_{\overrightarrow{n}},v_{\overrightarrow{n}}\}$ form a bipartition of $w_{\overrightarrow{n}}$ and
(7) $d_{R}(w)$ $=d_{R}(w_{\overleftarrow{n}})+d_{R}(w_{\overrightarrow{n}}),$
(8) $d_{R}(u)$ $=d_{R}(u_{\overleftarrow{n}})+d_{R}(u_{\overrightarrow{n}}),$
(9) $d_{R}(v)$ $=d_{R}(v_{\overleftarrow{n}})+d_{R}(v_{\overrightarrow{n}}).$
Note that $w_{\overleftarrow{n}}$, $u_{\overleftarrow{n}}$, $v_{\overleftarrow{n}}$ and $w_{\overrightarrow{n}}$, $u_{\overrightarrow{n}}$, $v_{\overrightarrow{n}}$ belong to different types of Coxeter groups: the left words are of type $B$ and the right words are of type $A$.
@@
Here, $w=4\bar{5}\bar{2}\bar{6}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}31|\bar{1}\bar{3}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}625\bar{4}}$ with $u=4\bar{5}\bar{2}\bar{6}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}3\bar{1}|1\bar{3}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}625\bar{4}}$ and $v=\bar{6}\bar{5}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bar{4}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bar{3}}\bar{2}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1}|{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bar{1}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}3}4{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}56}$ forming a bipartition of $w$. We obtain that $u_{\overleftarrow{n}}=2\bar{1}|1\bar{2}$ and $v_{\overleftarrow{n}}=\bar{2}1|\bar{1}2$ form a bipartition of $w_{L}=21|\bar{1}\bar{2}$ in $W_{2}$, while $u_{\overrightarrow{n}}=4231$ and $v_{\overrightarrow{n}}=1234$ form a (trivial) bipartition of $w_{\overrightarrow{n}}=u_{\overrightarrow{n}}$ in $\mathcal{S}_{3}$. Moreover, we have
$4$ $=$ $d_{R}(w)$ $=$ $d_{R}(4\bar{5}\bar{2}\bar{6}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}31|.\bar{1}.\bar{3}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6.25.\bar{4}})$ $=$ $d_{R}(21|.\bar{1}.\bar{2})+d_{R}(4.23.1)$ $=$ $2+2$
$3$ $=$ $d_{R}(u)$ $=$ $d_{R}(4\bar{5}\bar{2}\bar{6}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}3\bar{1}|1.\bar{3}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6.25.\bar{4}})$ $=$ $d_{R}(2\bar{1}|1.\bar{2})+d_{R}(4.23.1)$ $=$ $1+2$
$1$ $=$ $d_{R}(v)$ $=$ $d_{R}(\bar{6}\bar{5}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bar{4}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bar{3}}\bar{2}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1|.\bar{1}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}3}4{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}56})$ $=$ $d_{R}(\bar{2}1|.\bar{1}2)+d_{R}(1234)$ $=$ $1+0$
@@
As in type $A$, the proof is by induction on $n\in\mathbb{N}^{*}$. The result is true for $n=1$ since $W_{1}$ is of type $B_{1}=A_{1}$. Now, let $\{u,v\}$ be a bipartition of an element $w\in W_{n}$ with $n\geq 2$. We suppose that the statement of Conjecture 1 is satisfied for all $w\in W_{k}$ with $1\leq k<n$ such that $n$ is in positive position in $w$.
Suppose first that $|{\overleftarrow{n}}(w)|=k$ with $k>0$. By Proposition 4.18, $\{u_{\overleftarrow{n}},v_{\overleftarrow{n}}\}$ is a bipartition of $w_{\overleftarrow{n}}$, and $\{u_{\overrightarrow{n}},v_{\overrightarrow{n}}\}$ is a bipartition of $w_{\overrightarrow{n}}$, $w_{\overleftarrow{n}}$ is an element of $W_{k}$, with $k<n$, and $w_{\overrightarrow{n}}$ is an element of $A_{n-1-k}$. Therefore $\{u_{\overleftarrow{n}},v_{\overleftarrow{n}}\}$ satisfies Conjecture 1 by induction while $\{u_{\overrightarrow{n}},v_{\overrightarrow{n}}\}$ satisfies Conjecture 1 using Theorem 3.1 in the case of type $A$. Hence:
[…]
@@
Now suppose that ${\overleftarrow{n}}(w)=\emptyset$. Therefore $n$ is positive position 11 (as in Example 4.16). Hence $w_{\overrightarrow{n}}\not=w$ is a standardization of the letters in positive positions in $w$ and an element of $\mathcal{S}_{n-1}$ such that $d_{R}(w_{\overrightarrow{n}})=d_{R}(w)$. By Proposition 4.18, $\{u_{\overrightarrow{n}},v_{\overrightarrow{n}}\}$ form a bipartition of $w_{\overrightarrow{n}}$ with $d_{R}(u_{R})=d_{R}(u)$, $d_{R}(v_{\overrightarrow{n}})=d_{R}(w)$. Applying Theorem 3.1 in the case of type type $A$, we obtain $d_{R}(w)=d_{R}(u)+d_{R}(v)$. ∎
@@
Let $w\in W_{n}$ such that $n$ is in positive position in $w$. Let $\{u,v\}$ be a bipartition of an element $w$. Then there is no inversion $(a,c)$ of $u$ or $v$ such that $a\in\mathcal{F}(w)\cup{\overleftarrow{n}}(w)$ and $c\in{\overrightarrow{n}}(w)$, nor such that $a\in\mathcal{F}(w)$ and $c\in{\overleftarrow{n}}(w)\cup{\overrightarrow{n}}(w)$. In particular, $n$ is placed after all letters of $\mathcal{F}(w)$ and ${\overleftarrow{n}}(w)$ in both $u$ and $v$.
@@
Let $w\in W_{n}$ such that $n$ is in positive position in $w$. Let $\{u,v\}$ be a bipartition of an element $w$. If there is a descent in the word $u$ (or in the word $v$) between two letters $c$ and $a$, then either $a,c\in{\overrightarrow{n}}(w)$, or $a,c\in{\overleftarrow{n}}(w)$, or $a,c\in\mathcal{F}(w)$. In other words, a descent can only happen between two letters of similar colors.
@@
Let $w\in W_{n}$ such that $n$ is in positive position in $w$. Let $\{u,v\}$ be a bipartition of an element $w$. Then if there is a descent in $u$ (or in $v$) between two letters $c$ and $a$, then either $a,c\in{\overrightarrow{n}}(w)$ or $a,c\in{\overleftarrow{n}}(w)\cup\mathcal{F}(w)$.
@@
Consider the form of the word $w$ in Lemma 4.17 and observe that the definition of the left letters in type $A$ (everything before the letter $n$ in $w$) corresponds to ${\overleftarrow{n}}(w)\cup\mathcal{F}(w)$. Similarly, the right letters of type $A$ correspond to ${\overrightarrow{n}}(w)$ in type $B$. The exact same line of reasoning used to prove Proposition 3.19 applies here with Lemma 4.20 and Proposition 4.5. ∎
@@
Suppose that $u$ is of the form $\dots ca\dots$ with $a<c$. Note that we do not specify if the descent is in positive, negative or 0 position as this does not impact the proof. In particular, we could have $a=\bar{c}$ if the descent is in position 0. Lemma 4.23 states that either $a,c\in{\overrightarrow{n}}(w)$ or $a,c\in{\overleftarrow{n}}(w)\cup\mathcal{F}(w)$. By Lemma 4.20, we cannot have $c\in{\overleftarrow{n}}(w)$ and $a\in\mathcal{F}(w)$. The only case to consider is then $c\in\mathcal{F}(w)$ and $a\in{\overleftarrow{n}}(w)$ which we suppose by contradiction. As ${\overleftarrow{n}}(w)$ is symmetric, we also have $\bar{a}\in{\overleftarrow{n}}(w)$ and in particular $\bar{a}\neq c$ and the descent is not in position 0. By symmetry, this implies that there is a descent between $\bar{a}\in{\overleftarrow{n}}(w)$ and $\bar{c}\in{\overrightarrow{n}}(w)$ (because $\mathcal{F}(w)=\overline{{\overrightarrow{n}}(w)}$) which is impossible accordingly to Lemma 4.23. ∎
@@
Let $w\in W_{n}$ such that $n$ is in positive position in $w$. Let $\{u,v\}$ be a bipartition of an element $w$. Assume that $u$ (or $v$) is of the form $\dots b\dots c\dots$ with $b$ and $c$ both in the same set $I$ such that: (1) $I={\overleftarrow{n}}(w)$, $I={\overrightarrow{n}}(w)$ or $I=\mathcal{F}(w)$; (2) the subword between $b$ and $c$ is non-empty all its letters are in $\bar{n},n]\!]^{*}\setminus I$. Then $b<c$.
@@
We suppose by contradiction that $u=\dots cu^{\prime}b\dots$ with $b<c$ and $b,c\in I$ with $I={\overleftarrow{n}}(w)$, $I={\overrightarrow{n}}(w)$ or $I=\mathcal{F}(w)$ such that $u^{\prime}$ is not empty and with letters not in $I$. We proceed to a case by case analysis and show that we always reach a contradiction
Case $b,c\in{\overleftarrow{n}}(w)$. The word $u^{\prime}$ is non-empty by hypothesis, we call $a$ the last letter of $u^{\prime}$ so $u$ is of the form $\dots c\dots ab\dots$. As $a\notin{\overleftarrow{n}}(w)$, there is no descent between $a$ and $b$ by Proposition 4.21 so $a<b$. In particular, $(a,c)\in\operatorname{Inv}(u)$. As $c\in{\overleftarrow{n}}(w)$, by Lemma 4.20, we cannot have $a\in\mathcal{F}(w)$, so $a\in{\overrightarrow{n}}(w)$. Besides, as $(a,c)\in\operatorname{Inv}(u)$, we have $(a,c)\not\in\operatorname{Inv}(v)$. By Lemma 4.20, we also know that $n$ is placed after all letters of ${\overleftarrow{n}}(w)$ in both $u$ and $v$, this gives
@@
We see that $(a,n)\not\in\operatorname{Inv}(u)$ and $(a,n)\not\in\operatorname{Inv}(v)$. This contradicts the fact that $(a,n)\in\operatorname{Inv}(w)$ since $a$ is a right letter of $w$.
Case $b,c\in{\overrightarrow{n}}(w)$. We call $a$ the last letter of $u^{\prime}$ so that $u=\dots c\dots ab\dots$. As before, because of Proposition 4.21, $a<b$. We have $a\in{\overleftarrow{n}}(w)\cup\mathcal{F}(w)$ and $(a,c)\in\operatorname{Inv}(u)$: this contradicts Lemma 4.20.
Case $b,c\in\mathcal{F}(w)$. Obtained by symmetry of the case $b,c\in{\overrightarrow{n}}(w)$. ∎
@@
Let $w\in W_{n}$ such that $n$ is in positive position in $w$. Let $\{u,v\}$ be a bipartition of an element $w$. By a similar argument as in the case of type $A$, we obtain that $\{u_{\overleftarrow{n}},v_{\overleftarrow{n}}\}$ form a bipartition of $w_{\overleftarrow{n}}$ and $\{u_{\overrightarrow{n}},v_{\overrightarrow{n}}\}$ form a bipartition of $w_{\overrightarrow{n}}$, by simply restricting the sets of inversions to the left word (resp. right word) letters and using Lemma 4.4.
We first show Equality (7). We use Lemma 4.17 and suppose that $n$ is in position $k>0$. There is a descent at position $0\leq i<k-1$ in $w$ if and only if there is a descent at position $i$ in $w_{\overleftarrow{n}}$ as $w_{\overleftarrow{n}}$ is the type $B$ standardization of the middle factor of $w$. Similarly, there is a descent at position $i>k$ in $w$ if and only if there is a descent at position ik+1i-k+1 in $w_{\overrightarrow{n}}$. Besides, there is no descent in $w$ at position k1k-1. We find $d_{R}(w)=d_{R}(w_{\overleftarrow{n}})+d_{R}(w_{\overrightarrow{n}})$ (as $d_{R}(w)$ and $d_{R}(w_{\overleftarrow{n}})$ only count descents in nonnegative positions).
@@
$d_{R}(u)\leq d_{R}(u_{\overleftarrow{n}})+d_{R}(u_{\overrightarrow{n}}).$
Indeed, suppose that we have a descent $ca$ in $u$ in positive position. Either $a,c\in{\overleftarrow{n}}(w)$ and the descent is kept in $u_{\overleftarrow{n}}$, still in positive position. Or $a,c\in{\overrightarrow{n}}(w)$ and the descent is kept in $u_{\overrightarrow{n}}$. Finally, if $a,c\in\mathcal{F}(w)$, then $u$ also has a descent in negative position between $\bar{a}$ and $\bar{c}$ in ${\overrightarrow{n}}(w)$ which was not counted in $d_{R}(u)$ but is kept in $u_{\overrightarrow{n}}$ and counted in $d_{R}(u_{\overrightarrow{n}})$.